OpenStudy (anonymous):

Hey can you help me solve this? -2v-7=-23

OpenStudy (solomonzelman):

multiply everything in the equation times -1. that will give you \[2v+7=23\] \[2v+7\color{blue}{ -7 } =23\color{blue}{ -7 } \]\[2v=16\]\[2v\color{blue}{ \div 2 } =16\color{blue}{ \div 2 }\]\[\color{red}{ v~=~? }\]

OpenStudy (anonymous):

-2V-7=23 +7 +7 -2V=30 divide by 2 -v=15

OpenStudy (the_fizicx99):

^ So wrong lol.

OpenStudy (solomonzelman):

@suckerofmath re-read the problem.